# DreamControl: Control-Based Text-to-3D Generation with 3D Self-Prior

### 11 Dec 2023 v1

@@ 2 Related Work
Controllable Generation. Text input is a flexible control in generation tasks, while other conditions like image [27, 34], video [39, 3], and 3D sketch [19] are also available for guiding generation. ControlNet [38] supports text-to-image synthesis with additional conditions, e.g., edge, normal, etc. It allows 3D generators to create 3D content with the guidance of 2D sketch [4], depth [37], and even video [29]. In this work, we use 2D conditions to maintain geometry consistency, which are rendered from our 3D prior.
@@ 3 Preliminaries
$\nabla_{\theta}\mathcal{L}_{\text{SDS}}(\theta)=\mathbb{E}_{t,\epsilon}\left[\omega(t)\left(\hat{\epsilon}_{\phi}(x_{t},t,y)-\epsilon\right)\frac{\partial x}{\partial\theta}\right],$ (1)
where $\hat{\epsilon}_{\phi}$ is the noise predicted by $\phi$.
@@
$\mathbb{E}_{t,\bm{\epsilon}}\left[\omega(t)\left(\hat{\epsilon}_{\phi}(x_{t},t,y)-\hat{\epsilon}_{\theta}(x_{t},t,c,y)\right)\frac{\partial x}{\partial\theta}\right],$ (2)
where $\hat{\epsilon}_{\theta}(x_{t},t,c,y)$ is the noise predicted by rendered images. In practice, it can be regarded as LoRA [8] (Low-Rank Adaption) conditioned with camera poses $c$, which is supervised by a standard diffusion loss as,
[…]
@@ 4 DreamControl
$\mathop{\text{min}}\limits_{\theta\in\Theta}\mathcal{L}_{\text{SDS}}(\theta)\iff\mathop{\text{min}}\limits_{\theta\in\Theta}D_{\text{KL}}(q^{\theta}_{t}(x_{t}|y,c)||p_{t}(x_{t}|y,c)),$ (4)
where $q^{\theta}_{t}$ and $p_{t}$ the distribution probability of the NeRF scene $\theta$ and a pre-trained 2D diffusion model, respectively. However, there exists a viewpoint bias in the 2D model, which means it may not be able to acquire an accurate distribution $p_{t}(x_{t}|y,c)$ for a given camera pose $c$. When overfitting to the biased distribution $p_{t}$, optimization target in Eq. 4 can degenerate to $D_{\text{KL}}(q^{\theta}_{t}(x_{t}|y)||p_{t}(x_{t}|y,\widetilde{c}))$, where $\widetilde{c}$ is the most frequent viewpoint. As a result, generated results may look similar in any camera view, giving rise to the Janus problem.
@@ 4.2 Control-Based Score Distillation
Specifically, for one optimization step, we render an RGB image $x$ from $\theta$ and a conditional image $\hat{x}$ from $\hat{\theta}$ in the same camera pose. Then, ControlNet can supervise the generation by the predicted noise $\hat{\epsilon}_{\phi}(x_{t},t,\hat{x}_{t},y)$ according to Eq. 1, which is demonstrated effective in previous works [29, 37]. However, these methods generally require high-quality conditions for the generation in ControlNet, e.g., depth maps or DSLR photos. Strong conditions such as depth may restrict the diversity of generated content. Meanwhile, our 3D prior $\hat{\theta}$ can only exhibit poor texture information, making it hard to extract photo-realistic images for generative control.
As for the issue of diversity, we take VSD [32] into consideration, which expands the generation range of a given prompt. Since photo-realistic images are not available from our 3D priors, we use the boundary mask from $\hat{\theta}$ as the condition $\hat{x}$, further weakening the control restriction. However, simply replacing the pre-trained diffusion model with ControlNet doesn’t work well, as the pre-trained term $\hat{\epsilon}_{\phi}$ and the LoRA term $\hat{\epsilon}_{\theta}$ vary a lot, making it hard for the convergence of Eq. 3. To stabilize the optimization, we propose a conditional LoRA and a weighted score.
Conditional LoRA. In VSD, a LoRA is adopted to predict the noise of current NeRF scene $\hat{\epsilon}_{\theta}(x_{t},t,c,y)$, in which a camera pose $c$ is additionally embedded. We are concerned that camera pose is a high-level semantic concept, which is hard to tokenize and imbibe through LoRA. The LoRA term $\hat{\epsilon}_{\theta}$ may predict a noise unrelated to $c$, enlarging the gap with the pre-trained term $\hat{\epsilon}_{\phi}$. Instead, we replace camera pose $c$ with the normal map $x_{t}^{n}$ rendered by itself. Accordingly, a lightweight control LoRA is adopted to predict noise conditioned by $x_{t}^{n}$. The training objective is similar to Eq. 3.
Weighted Score. The prediction of the LoRA term $\hat{\epsilon}_{\theta}$ in the early stage lacks practical significance, as NeRF has not yet generated meaningful objects. To mitigate the possible disruption caused by LoRA at those early steps, we propose to incorporate a coefficient $\lambda$ into its loss term, which is changed along with training steps. Specifically, we rewrite $\hat{\epsilon}_{\phi}(x_{t},t,y)-\hat{\epsilon}_{\theta}(x_{t},t,c,y)$ in Eq. 2 as two terms,
$(\hat{\epsilon}_{\phi}(x_{t},t,\hat{x}_{t},y)-\epsilon)-\lambda(\hat{\epsilon}_{\theta}(x_{t},t,x_{t}^{n},y)-\epsilon).$ (7)
@@ 4.3 Implementation Details
We implement DreamControl based on threestudio [6]. In the first stage, we adopt DeepFloyd [31] guidance in the SDS optimization. In the second stage, we use the scribble version of ControlNet [38] v1.1 as the pre-trained term and the stable diffusion [26] v1.5 injected with ContraolLoRA [7] as the LoRA term. 3D prior is the rendered density mask from the NeRF result in the first stage, further processed by HEDdetector [35] implemented in ControlNet. The LoRA condition is the rendered normal map from the current NeRF. The classifier-free guidance scale (CFG) is set as 7.5 and 1.0 in pre-trained and LoRA terms, respectively. We increase the loss coefficient $\lambda$ from 0.5 to 1.0 linearly in the first 5,000 iterations. Please refer to the Suppl. for more details.

### 12 Mar 2024 v2

@@ 2 Related Work
Controllable Generation. Text input is a flexible control in 3D tasks [9], while other conditions like image [29, 37], video [42, 3], and 3D sketch [20] are also available for guiding generation. ControlNet [41] supports text-to-image synthesis with additional conditions, e.g., edge, normal, etc. It allows 3D generators to create 3D content with the guidance of 2D sketch [4], depth [40], and even video [31]. In this work, we use 2D conditions to maintain geometry consistency, which are rendered from our 3D prior.
@@ 3 Preliminaries
$\nabla_{\theta}\mathcal{L}_{\text{SDS}}(\theta)=\mathbb{E}_{t,\bm{\epsilon}}\left[\omega(t)\left(\hat{\bm{\epsilon}}_{\phi}(\bm{x}_{t},t,y)-\bm{\epsilon}\right)\frac{\partial\bm{x}}{\partial\theta}\right],$ (1)
where $\hat{\bm{\epsilon}}_{\phi}$ is the noise predicted by $\phi$.
@@
$\mathbb{E}_{t,\bm{\bm{\epsilon}}}\left[\omega(t)\left(\hat{\bm{\epsilon}}_{\phi}(\bm{x}_{t},t,y)-\hat{\bm{\epsilon}}_{\theta}(\bm{x}_{t},t,c,y)\right)\frac{\partial\bm{x}}{\partial\theta}\right],$ (2)
where $\hat{\bm{\epsilon}}_{\theta}(\bm{x}_{t},t,c,y)$ is the noise predicted by rendered images. In practice, it can be regarded as LoRA [8] (Low-Rank Adaption) conditioned with camera poses $c$, which is supervised by a standard diffusion loss as,
$\mathbb{E}_{t\sim\mathcal{U}(0,1),\bm{\epsilon}\sim\mathcal{N}(0,\mathbf{\textit{I}})}\left[\|\hat{\bm{\epsilon}}_{\theta}(\alpha_{t}\bm{x}_{t}+\sigma_{t}\bm{\epsilon},t,c,y)-\bm{\epsilon}\|_{2}^{2}\right].$ (3)
@@ 4 DreamControl
$\mathop{\text{min}}\limits_{\theta\in\Theta}\mathcal{L}_{\text{SDS}}(\theta)\iff\mathop{\text{min}}\limits_{\theta\in\Theta}D_{\text{KL}}(q^{\theta}_{t}(\bm{x}_{t}|y,c)||p_{t}(\bm{x}_{t}|y,c)),$ (4)
where $q^{\theta}_{t}$ and $p_{t}$ the distribution probability of the NeRF scene $\theta$ and a pre-trained 2D diffusion model, respectively. However, there exists a viewpoint bias in the 2D model, which means it may not be able to acquire an accurate distribution $p_{t}(\bm{x}_{t}|y,c)$ for a given camera pose $c$. When overfitting to the biased distribution $p_{t}$, optimization target in Eq. 4 can degenerate to $D_{\text{KL}}(q^{\theta}_{t}(\bm{x}_{t}|y)||p_{t}(\bm{x}_{t}|y,\widetilde{c}))$, where $\widetilde{c}$ is the most frequent viewpoint. As a result, generated results may look similar in any camera view, giving rise to the Janus problem.
@@ 4.2 Control-Based Score Distillation
Specifically, for one optimization step, we render an RGB image $\bm{x}$ from $\theta$ and a conditional image $\hat{\bm{x}}$ from $\hat{\theta}$ in the same camera pose. Then, ControlNet can supervise the generation by the predicted noise $\hat{\bm{\epsilon}}_{\phi}(\bm{x}_{t},t,\hat{\bm{x}}_{t},y)$ according to Eq. 1, which is demonstrated effective in previous works [31, 40]. However, these methods generally require high-quality conditions for the generation in ControlNet, e.g., depth maps or DSLR photos. Strong conditions such as depth may restrict the diversity of generated content. Meanwhile, our 3D prior $\hat{\theta}$ can only exhibit poor texture information, making it hard to extract photo-realistic images for generative control.
As for the issue of diversity, we take VSD [35] into consideration, which expands the generation range of a given prompt. Since photo-realistic images are not available from our 3D priors, we use the boundary mask from $\hat{\theta}$ as the condition $\hat{x}$, further weakening the control restriction. However, simply replacing the pre-trained diffusion model with ControlNet doesn’t work well, as the pre-trained term $\hat{\bm{\epsilon}}_{\phi}$ and the LoRA term $\hat{\bm{\epsilon}}_{\theta}$ vary a lot, making it hard for the convergence of Eq. 3. To stabilize the optimization, we propose a conditional LoRA and a weighted score.
Conditional LoRA. In VSD, a LoRA is adopted to predict the noise of current NeRF scene $\hat{\bm{\epsilon}}_{\theta}(\bm{x}_{t},t,c,y)$, in which a camera pose $c$ is additionally embedded. We are concerned that camera pose is a high-level semantic concept, which is hard to tokenize and imbibe through LoRA. The LoRA term $\hat{\bm{\epsilon}}_{\theta}$ may predict a noise unrelated to $c$, enlarging the gap with the pre-trained term $\hat{\bm{\epsilon}}_{\phi}$. Instead, we replace camera pose $c$ with the normal map $\bm{x}_{t}^{n}$ rendered by itself. Accordingly, a lightweight control LoRA is adopted to predict noise conditioned by $\bm{x}_{t}^{n}$. The training objective is similar to Eq. 3.
Weighted Score. The prediction of the LoRA term $\hat{\bm{\epsilon}}_{\theta}$ in the early stage lacks practical significance, as NeRF has not yet generated meaningful objects. To mitigate the possible disruption caused by LoRA at those early steps, we propose to incorporate a coefficient $\lambda$ into its loss term, which is changed along with training steps. Specifically, we rewrite $\hat{\bm{\epsilon}}_{\phi}(\bm{x}_{t},t,y)-\hat{\bm{\epsilon}}_{\theta}(\bm{x}_{t},t,c,y)$ in Eq. 2 as two terms,
$(\hat{\bm{\epsilon}}_{\phi}(\bm{x}_{t},t,\hat{\bm{x}}_{t},y)-\bm{\epsilon})-\lambda(\hat{\bm{\epsilon}}_{\theta}(\bm{x}_{t},t,\bm{x}_{t}^{n},y)-\bm{\epsilon}).$ (7)
@@ 4.3 Implementation Details
We implement DreamControl based on threestudio [6]. In the first stage, we adopt DeepFloyd [34] guidance in the SDS optimization. In the second stage, we use the scribble version of ControlNet [41] v1.1 as the pre-trained term and the stable diffusion [28] v1.5 injected with ContraolLoRA [7] as the LoRA term. 3D prior is the rendered density mask from the NeRF result in the first stage, further processed by HEDdetector [38] implemented in ControlNet. The LoRA condition is the rendered normal map from the current NeRF. The classifier-free guidance scale (CFG) is set as 7.5 and 1.0 in pre-trained and LoRA terms, respectively. We increase the loss coefficient $\lambda$ from 0.5 to 0.75 linearly in the first 5,000 iterations. Please refer to the Suppl. for more details.
@@ 6 Conclusion
Acknowledgements
This work was supported by National Key RD Program of China under Grant No. 2021ZD0112100, and the National Natural Science Foundation of China (NSFC) under Grant No. U19A2073.
Supplementary Material
In this supplementary, we provide more information on our implementation details (Section 7) and evaluation metrics (Section 8). To demonstrate the effectiveness of our control-based guidance on maintaining 3D prior, we provide a comparison with a two-stage ProlificDreamer (Section 9). For more visualization results in the format of $360^{\circ}$, please refer to the attached HTML file in results folder.
7 Implementation Details
[…]
Boundary Integrity Metric. Our terminated condition in the first stage is related to the density difference between all valid rays and edge rays in NeRF scenes. The edge is detected by HEDdetector [38] implemented in ControlNet. We calculate the average density of each ray set and terminate the optimization when $\Delta_{\mathbf{r}}$ is less than 0.1 for three consecutive checkpoints. We set a checkpoint for every 100 iterations. We also provide an example of generating corgi in Figure 12. As the density difference decreases, the NeRF representation gradually forms a solid geometry.
Mesh Extraction. Our two-stage optimization can achieve promising visual results under the representation of NeRF, while the surfaces may exhibit severe irregularities if we export NeRF results to mesh objects for applications like animation. To obtain higher-quality 3D models for broader applications, we further refine the mesh representation based on DMTet [32]. We employ the same ControlNet guidance under NeRF representation to optimize the refinement of SDF representation. Specifically, we convert NeRF into a DMTet object and retain its texture information. We initially optimize its geometry to achieve smooth and regular surfaces. Subsequently, we optimized its texture to remove the noise generated during the geometry optimization. Note that, the optimization guidance in the SDF phase is exactly the same with the NeRF phase. The parameters of conditional LoRA on the NeRF training are loaded before the SDF optimization, which is frozen during the geometry phase and unfrozen during the texture phase.
Figure 11: Visualization of generation in different view-dependent prompts. Generated corgis in the front view present a higher confidence than the other two views.
Figure 12: Visualization of the change in density difference $\Delta_{\mathbf{r}}$. As $\Delta_{\mathbf{r}}$ decreases, the NeRF representation gradually forms a solid geometry.
Inference Time. In the the first stage, as the optimization is terminated automatically based on our metric, it usually takes less than 10 minutes for the generation of 3D self-prior. In the second stage, we optimize the generation for 15,000 iterations, which spend around 2 hours for a given prompt input. The total inference time is basically equivalent to previous text-to-3D methods like Fantasia3D [2] and ProlificDreamer [35]. For further extracting the mesh representation, since the color information from the NeRF phase is already of high quality, in the SDF phase, we only need to optimize the geometry and suppress texture noise appropriately. As a result, the entire process of mesh extraction is relatively fast, taking around 30 minutes.
Pseudocode. We provide pseudocode in Algorithm 1,2 to summarize the NeRF phase and the SDF phase in our framework, respectively.
Algorithm 1 DreamControl - NeRF Phase
1:  Input: text prompt y
2:  Load: stable diffusion v1.5 $\phi_{\text{sd}}$ and DeepFloyd XL-v1.0 $\phi_{\text{if}}$
[…]
5:  for $t=10,20,\dots,100$ do
6:   Generate images $\bm{x}_{y_{1}}^{t}$, $\bm{x}_{y_{2}}^{t}$, $\bm{x}_{y_{3}}^{t}$ with $\phi_{\text{if}}$ in the timestep $t$
7:  end for
8:  $p^{\ast}_{1,2,3}=\text{softmax}(s_{\text{CLIP}}(y_{1,2,3},[\{\bm{x}_{y_{1}}^{t}\},\{\bm{x}_{y_{2}}^{t}\},\{\bm{x}_{y_{3}}^{t}\}]))$
[…]
10:  initialize a NeRF scene $\theta$
11:  while Not $\Delta_{\mathbf{r}}<0.1$ for three consecutive checkpoints do
12:   Sample a camera $c$ based on $p^{\ast}$ and a timestep $t$
13:   Render $\theta$ at pose $c$, for RGB image $\bm{x}=\mathbf{g}(\theta,c)$
14:   $\theta\leftarrow\theta-\mathbb{E}\left[\omega(t)\left(\hat{\bm{\epsilon}}_{\phi_{\text{if}}}(\bm{x}_{t},t,y)-\bm{\epsilon}\right)\frac{\partial\bm{x}}{\partial\theta}\right]$
15:  end while
16:  3D self-prior $\hat{\theta}=\theta$
[…]
18:  initialize a NeRF scene $\theta$ and a conditional LoRA $\phi_{\theta}$
19:  while not converged do
20:   Sample a camera $c$ and a timestep $t$
21:   Render $\theta$ for RGB image $\bm{x}$ and normal map $\bm{x}^{n}$ at pose $c$
22:   Render $\hat{\theta}$ and detect an edge mask $\hat{\bm{x}}$ at pose $c$
23:   Update $\lambda$, weighted score $\mathcal{L}_{\text{score}}(\bm{x}_{t},\hat{\bm{x}}_{t},\bm{x}_{t}^{n},t)=(\hat{\bm{\epsilon}}_{\phi_{\text{sd}}}(\bm{x}_{t},t,\hat{\bm{x}}_{t},y)-\bm{\epsilon})-\lambda(\hat{\bm{\epsilon}}_{\theta}(\bm{x}_{t},t,\bm{x}_{t}^{n},y)-\bm{\epsilon})$
24:   $\theta\leftarrow\theta-\mathbb{E}\left[\omega(t)\mathcal{L}_{\text{score}}(\bm{x}_{t},\hat{\bm{x}}_{t},\bm{x}_{t}^{n},t)\frac{\partial\bm{x}}{\partial\theta}\right]$
25:   $\phi_{\theta}\leftarrow\phi_{\theta}-\nabla_{\phi_{\theta}}\mathbb{E}\left[\|\hat{\bm{\epsilon}}_{\theta}(\alpha_{t}\bm{x}_{t}+\sigma_{t}\bm{\epsilon},t,c,y)-\bm{\epsilon}\|_{2}^{2}\right].$
26:  end while
27:  return  the optimized NeRF representation $\theta$
Algorithm 2 DreamControl - SDF Phase
1:  Input: text prompt y
2:  Load: $\phi_{\text{sd}}$, $\phi_{\theta}$, and $\theta$
[…]
4:  initialize a DMTet mesh from $\theta$, parameterized by $m$
5:  while not converged do
6:   Sample a camera $c$ and a timestep $t$
7:   Render $m$ for RGB image $\bm{x}$ and normal map $\bm{x}^{n}$ at pose $c$
8:   Render $\hat{\theta}$ and detect an edge mask $\hat{\bm{x}}$ at pose $c$
9:   $m\leftarrow m-\mathbb{E}\left[\omega(t)\left(\hat{\bm{\epsilon}}_{\phi_{sd}}(\bm{x}_{t},t,y)-\hat{\bm{\epsilon}}_{\theta}(\bm{x}_{t},t,\bm{x}_{t}^{n},y)\right)\frac{\partial\bm{x}}{\partial m}\right]$
10:  end while
[…]
12:  while not converged do
13:   Sample a camera $c$ and a timestep $t$
14:   Render $m$ for RGB image $\bm{x}$ and normal map $\bm{x}^{n}$ at pose $c$
15:   Render $\hat{\theta}$ and detect an edge mask $\hat{\bm{x}}$ at pose $c$
16:   $m\leftarrow m-\mathbb{E}\left[\omega(t)\left(\hat{\bm{\epsilon}}_{\phi_{sd}}(\bm{x}_{t},t,y)-\hat{\bm{\epsilon}}_{\theta}(\bm{x}_{t},t,\bm{x}_{t}^{n},y)\right)\frac{\partial\bm{x}}{\partial m}\right]$
17:   $\phi_{\theta}\leftarrow\phi_{\theta}-\nabla_{\phi_{\theta}}\mathbb{E}\left[\|\hat{\bm{\epsilon}}_{\theta}(\alpha_{t}\bm{x}_{t}+\sigma_{t}\bm{\epsilon},t,c,y)-\bm{\epsilon}\|_{2}^{2}\right].$
18:  end while
19:  return  the optimized SDF representation $m$
8 Evaluation Metrics
Janus Rate. To evaluate the geometry consistency, we count the occurrence rate of the Janus problem (JR). As shown in Figure 13, we count generation as a Janus problem based on the following situations: (1) multi-face, multi-hand, multi-leg, or similar issues; (2) obvious content drift; (3) serious paper-thin generation. We calculate JR as the number of inconsistent content out of the number of all generated objects.
Pick-Score. Pick-Score (PS) [13] is a CLIP-based scoring model, which is trained with Pick-a-Pic, a large, open dataset of text-to-image prompts and real users’ preferences over generated images. As a result, it can exhibit superhuman performance on the task of predicting human preferences. In 3D evaluation, we compare multi-view rendered images of generated objects to measure the preference. For a given text prompt $y$, we render N-view images $\{x_{y,i}^{k}\}_{i=1}^{N}$ from the generated object of method $k$. The pick score ps for $y$ is formulated as,
$\text{ps}(y)=\frac{1}{N}\sum_{i=1}^{N}s_{\text{pick}}(y,[\bm{x}_{y,i}^{1},\bm{x}_{y,i}^{2},\dots,\bm{x}_{y,i}^{k}]),$ (8)
where $s_{\text{pick}}$ is the CLIP-based scoring function. For each method $k$, the final pick score is calculated as the average score of all the text prompts, i.e., $\frac{1}{|Y|}\sum_{y\in Y}\text{ps}(y,k)$.
CLIP-Score. CLIP-Score (CS) is based on the CLIP [27] similarity. For each prompt $y$, we render one image from the corresponding generated object of method $k$, $\bm{x}^{k}_{y}$. The rendering viewpoint is a fixed camera pose at a $45^{\circ}$ angle of elevation and a $30^{\circ}$ angle of rotation. The final score is calculated as the average score of all the text prompts, i.e., $\frac{1}{|Y|}\sum_{y\in Y}s_{\text{CLIP}}(y,\bm{x}^{k}_{y})$. Note that, differently from Pick-Score, we don’t evaluate the CLIP-score on multi-view images because the Janus problem could ironically increase the CLIP similarity.
Figure 13: Visualization of Janus problems. (1) multi-face, multi-hand, multi-leg, or similar issues; (2) obvious content drift; (3) serious paper-thin generation.
9 Two-Stage Comparison
To further demonstrate that our control-based guidance can maintain the 3D self-prior obtained in the first stage, we compare our second stage with two-stage ProlificDreamer [35]. For the two-stage ProlificDreamer, we use the 3D self-prior in our first stage for initialization.
As shown in Figure 14, we take the prompt “A chimpanzee dressed like Henry VIII king of England” as an example. The final generated chimpanzee of ProlificDreamer is quite different from its initialization, presenting multiple faces and multiple hands. Although DreamControl only uses the edge condition, our result can successfully maintain the geometry of the input condition. The results demonstrate that continually optimizing a coarse shape may still lead to overfitting issues.
Figure 14: Visualization of two-stage generation comparison.
10 More Visualization Results
Please refer to the attached HTML file in results folder for more visualization results in the format of $360^{\circ}$ video.